# A Single Graph Convolution is All You Need: Efficient Grayscale Image Classification

### v4

@@ Abstract
Image classifiers often rely on convolutional neural networks (CNN) for their tasks, which are inherently more heavyweight than multilayer perceptrons (MLPs), which can be problematic in real-time applications. Additionally, many image classification models work on both RGB and grayscale datasets. Classifiers that operate solely on grayscale images are much less common. Grayscale image classification has diverse applications, including but not limited to medical image classification and synthetic aperture radar (SAR) automatic target recognition (ATR). Thus, we present a novel grayscale (single channel) image classification approach using a vectorized view of images. We exploit the lightweightness of MLPs by viewing images as a vector and reducing our problem setting to the grayscale image classification setting. We find that using a single graph convolutional layer batch-wise increases accuracy and reduces variance in the performance of our model. Moreover, we develop a customized accelerator on FPGA for the proposed model with several optimizations to improve its performance. Our experimental results on benchmark grayscale image datasets demonstrate the effectiveness of the proposed model, achieving vastly lower latency (up to 16$\times$ less) and competitive or leading performance compared to other state-of-the-art image classification models on various domain-specific grayscale image classification datasets.
Index Terms—  GCN, grayscale, image classification, MLP, low-latency
@@ 1 Introduction
A particular instance where low-latency machine learning is needed is grayscale image classification. For example, a targeting system on a satellite is costly, and decisions must be made using SAR efficiently and accurately. Examples like this are where low-latency grayscale image classification comes into play. It is often the case that image classifiers work on RGB datasets and grayscale image datasets, but seldom do modern image classifiers focus solely on the grayscale setting. RGB models are overkill for the grayscale setting, as the grayscale problem allows us to focus on a single channel. Models focusing on grayscale image classification are naturally more efficient, as they can concentrate on a single channel rather than three. Thus, many image classifiers that generalize to the grayscale image classification are not truly optimized for the grayscale case. For these reasons, we present a lightweight grayscale image classifier capable of achieving up to 16×16$\times$16 × lower latency than other state-of-the-art machine learning models.
From a trustworthy visual data processing perspective, the demand for grayscale image classification requires data to be collected from various domains with high resolution and correctness so that we can train a robust machine learning model. Additionally, recent advancements in machine learning rely on convolutional neural networks, which often suffer from high computation costs, large memory requirements, and many network parameters, resulting in poor inference latency, poor scalability, and weak trustworthiness.
The inherent novelties of our model are the following: our proposed method is the first that vectorizes an image in a fully connected manner and inputs the resultant into a single-layer graph convolutional network (GCN). We also find that a single GCN layer is enough to stabilize the performance of our shallow model. Additionally, our proposed method benefits from a batch-wise attention term, allowing our shallow model to capture interdependencies between images and form connections for classification. Finally, by focusing on grayscale imagery, we can focus on a streamlined method for grayscale image classification rather than concentrating on the RGB setting. A result of these novelties is extremely low latency and high throughput for image classification.
@@
  • Experiments show that our model achieves competitive or leading accuracy with respect to other popular state-of-the-art models while vastly reducing latency and model complexity.
  • We implement our model on a state-of-the-art FPGA board, Xilinx Alveo U200. Compared with state-of-the-art GPU implementation, our FPGA implementation achieves a 2.78×2.78$\times$2.78 × speedup in latency and a 2.1×2.1$\times$2.1 × speedup in throughput.
@@ 2 Problem Definition
Latency refers to the total time (from start to finish) it takes to gather predictions for a model in one batch (a standard approach). A lightweight machine learning model aims to maximize throughput and accuracy while minimizing latency.
@@ 4 Overview and Architecture
Our model vectorizes a batch of images, allowing us to use a fully connected layer pixel-wise for low computation time rather than relying on convolutional neural networks. The proposed method uses the resultant from FCs and constructs a graph batch-wise, where each node corresponds to the flattened image outputted from an FC. We then apply a batch-wise attention term, which is inputted into an FC for classification. [^1] We make our code publicly available at https://github.com/GECCOProject/GECCO
@@
Fully Connected Layer. We input $\mathbf{X}_{1}$ into a fully connected layer (FC) with output dimensionality $D_{out}$. Formally, $\mathbf{X}_{2}=\sigma(\mathbf{X}_{1}\mathbf{W}_{1}+\mathbf{b}_{1})$, where $\sigma$ is the ReLU function, $\mathbf{W}_{1}$ is a learned weight matrix, $\mathbf{b}_{1}$ is a bias term, and $\mathbf{X}_{2}\in\mathbb{R}^{B\times D_{out}}$.
@@
Applying graph convolution to $\mathbf{X}_{3}$ yields $\mathbf{X}_{4}$. Formally, $\mathbf{X}_{4}=\sigma\left(\mathbf{A}\mathbf{X}_{3}\mathbf{W}_{2}\right)$, where $\mathbf{W}_{2}$ is a learned weight matrix. After the graph convolution, we apply batch normalization and max-pooling operations to $\mathbf{X}_{4}$, resulting in a dimensionality of $\mathbb{R}^{B\times\lfloor D_{out}/2\rfloor}$.
@@
$\mathbf{X}_{5}$ $=\left(\frac{\sigma\left(\mathbf{X}_{4}\mathbf{X}^{\top}_{4}\right)}{\sum_{i=1}^{B}\sum_{j=1}^{B}\sigma\left(\mathbf{X}_{4}\mathbf{X}_{4}^{\top}\right)_{ij}}\right)\mathbf{X}_{4}$
@@
The residual connection is defined $\mathbf{X}_{6}=\mathbf{X}_{5}+\mathbf{X}_{4}$. The residual term makes the learning process easier and more stable. By multiplying a softmax-like term with the output of a previous graph convolution, we weigh the correspondence of each pixel compared to other similar pixels within similar images batch-wise. We then feed the residual term into an FC inputted into the softmax function for classification results.
@@ 5.1 Datasets
Datasets from several domains are examined to gauge the effectiveness of GECCO in diverse settings. We use the popular MNIST dataset [12], Fashion-MNIST [13], MSTAR, and CXR [14] summarized in the following manner:
  • MNIST is a grayscale handwritten dataset with $(28,28)$ pixel image sizes and $10$ different object categories. The training size for this dataset is $60000$, and the testing size is $10000$.
  • Fashion-MNIST contains $(28,28)$ sized grayscale images from $10$ categories with a training size of $60000$, and a testing size of $10000$.
@@ 5.2.1 Backbone
Table 1: Performance of GCN Layers Varying on MNIST
@@
GCN [15] $97.12\%$ $54.58\pm 8.47$ $4.68\pm 0.72$
TAGConv [16] $97.36\%$ $58.39\pm 9.01$ $4.37\pm 0.67$
SAGEConv [17] $97.72\%$ $58.84\pm 8.98$ $4.32\pm 0.66$
ChebConv [18] $97.79\%$ $51.98\pm 8.65$ $4.91\pm 0.82$
ResConv [19] $\bf{97.84\%}$ $\bf{62.44\pm 9.87}$ $\bf{4.10\pm 0.65}$
We use the following hyperparameters for our experiments.
@@ 5.2.2 Experimental Performance
In Tables 3, 4, 5, and 6, we present a summary of our findings. We report the best-performing accuracy, average throughputs, and latencies with their standard deviations. Our model outperforms every other model in terms of throughput and latency across all datasets, leads accuracy on the MSTAR dataset, and performs competitively in terms of accuracy on all datasets.
@@
Table 3: MNIST Performance
[…]
Swin-T $98.30\%$ $4.69\pm 0.21$ $54.53\pm 2.44$
SS-ViT $98.09\%$ $8.38\pm 1.42$ $30.46\pm 5.17$
VGG16 $\bf{99.54\%}$ $39.02\pm 4.98$ $6.55\pm 0.84$
FastViT $98.44\%$ $7.51\pm 0.51$ $34.05\pm 2.31$
ResNet34 $99.06\%$ $12.88\pm 1.02$ $19.86\pm 1.57$
GECCO $97.84\%$ $\bf{62.44\pm 9.87}$ $\bf{4.10\pm 0.65}$
Table 4: Fashion-MNIST Performance
[…]
Swin-T $88.73\%$ $4.58\pm 0.31$ $55.81\pm 3.78$
SS-ViT $87.88\%$ $8.27\pm 1.56$ $30.97\pm 5.83$
VGG16 $\bf{92.48\%}$ $24.98\pm 5.77$ $10.24\pm 2.36$
FastViT $87.94\%$ $7.28\pm 0.75$ $35.07\pm 3.63$
ResNet34 $88.50\%$ $11.98\pm 1.05$ $21.35\pm 1.87$
GECCO $87.87\%$ $\bf{59.08\pm 9.04}$ $\bf{4.32\pm 0.66}$
@@
Swin-T $86.04\%$ $5.44\pm 0.37$ $46.98\pm 3.20$
SS-ViT $95.61\%$ $9.14\pm 1.72$ $27.97\pm 5.26$
VGG16 $93.13\%$ $6.75\pm 1.34$ $37.89\pm 7.52$
FastViT $91.78\%$ $4.16\pm 0.52$ $61.44\pm 7.69$
ResNet34 $98.64\%$ $12.44\pm 0.84$ $20.48\pm 1.39$
GECCO $\bf{99.42\%}$ $\bf{67.17\pm 15.24}$ $\bf{3.80\pm 0.86}$
@@
Swin-T $73.66\%$ $1.08\pm 0.21$ $236.71\pm 46.09$
SS-ViT $71.09\%$ $4.09\pm 0.84$ $62.35\pm 12.85$
VGG16 $\bf{82.01\%}$ $3.03\pm 1.02$ $84.10\pm 28.43$
FastViT $75.46\%$ $4.24\pm 1.00$ $60.30\pm 14.24$
ResNet34 $78.31\%$ $2.41\pm 0.44$ $105.84\pm 19.39$
GECCO $79.87\%$ $\bf{14.04\pm 3.19}$ $\bf{18.05\pm 4.14}$
@@ 5.2.3 Model Complexity Metrics
Model complexity metrics for this paper include the number of multiply-accumulate operations (MACs), the number of model parameters, the model size, and the number of layers. In other words, suppose accumulator $a$ counts an operation of arbitrary $b,c\in\mathbb{R}$. We count the number of multiply-accumulate operations as $a\leftarrow a+(b\times c)$. Additionally, the layer count metric is an essential factor of latency. Increasing the number of layers will also improve the latency of a model’s inference. The goal of an effective machine learning model is to maximize throughput while minimizing the number of MACs and the number of layers, in our case.
@@
GECCO $\bf{5.89\times 10^{6}}$ $\bf{1.90\times 10^{4}}$ $\bf{0.075}$ $\bf{10}$
@@ 5.2.4 Ablation Study
We perform an ablation study to verify that the components of our proposed model contribute positively to the overall accuracy.
@@
Mini-batch GNN Weighted Sum Residual Term Accuracy on MNIST
$97.84\%$
$93.09\%$
$91.64\%$
$86.32\%$
@@
Fig. 2: Accuracy on MNIST vs. Number of Graph Convolutional Layers
@@ 5.3 Discussion
It is difficult for our model to generalize to the RGB setting. We attribute this challenge to the vectorization process that our model uses. Learning on three channels poses a complexity challenge, as GECCO is very shallow and simple, thus making it challenging to learn on three separate channels. Additionally, our model is optimized for a low-complexity dataset regime, as they are much too complex for our shallow model. Datasets like grayscale ImageNet or CIFAR variants are unsuitable for GECCO as they pose a complexity challenge.
Our proposed method does not make use of positional embeddings or class tokens. GECCO can learn essential features using the weighted residual term. Additionally, we tested the addition of positional embeddings and class tokens and found no improvement in accuracy across various datasets. We note that the $\mathbf{X}_{5}$ residual term adds positional awareness to the model.
@@ 5.4 FPGA Implementation
We develop an accelerator for the proposed model on a state-of-the-art FPGA, Xilinx Alveo U200 board, to further highlight the model’s efficiency and compatibility with hardware. It has 3 Super Logic Regions (SLRs), 4 DDR memory banks, 1182k Look-up tables, 6840 DSPs, 75.9 Mb of BRAM, and 270 Mb of URAM. The FPGA kernels are developed using the Xilinx High-level Synthesis (HLS) tool to expedite the design process.
@@
We perform the following optimizations for our FPGA design: (1) Portable design: We design a parameterized hardware template using HLS. It is portable to different FPGA platforms, including embedded and data-center FPGAs. (2) Resource sharing: The model is executed layer-by-layer. Each layer in the model is decomposed into basic kernel functions. The basic kernel functions, including matrix multiplication, elementwise activation, column-wise and row-wise summations, max pooling, and various other elementwise operations, are implemented separately and subsequently invoked within their corresponding layers. Due to the reuse of these fundamental kernel functions across multiple layers, FPGA resources are shared among the different layers, maximizing runtime hardware resource utilization. (3) One-load strategy: We employ a one-time data load strategy to load the required data from DDR only once. All other data required for the computations are stored in on-chip memory, reducing inference latency. Figure 3 illustrates the overall hardware architecture of our design.
@@
We use the Vitis Analyzer tool to provide insights into resource utilization, latency, and throughput. Table 9 reports the results obtained for the MNIST dataset. Given the model’s compact design and resource efficiency, it can be accommodated within a single SLR. Hence, we deploy multiple accelerator instances across multiple SLRs, each with one instance. This increases the inference throughput. In table 9 shows the latency obtained for a single inference and the throughput achieved by running the design on 3 SLRs concurrently.
@@
We compare our FPGA implementation with the baseline GPU implementation. The GPU baseline is executed on an NVIDIA RTX A5000 GPU, which operates at 1170 MHz and has a memory bandwidth of 768 GB/s. On the other hand, the FPGA operates at 300 MHz and has an external memory bandwidth of 77 GB/s. The GPU baseline is comparable with the FPGA in terms of the platform. Although the GPU has higher peak performance and memory bandwidth, our FPGA implementation achieves a latency reduction of 2.78×2.78$\times$2.78 × and a throughput improvement of 2.1×2.1$\times$2.1 ×. This speedup is attributed to the optimizations mentioned above that were adopted in our implementation.
@@
Table 9: Resource Utilization (per SLR), Latency, and Throughput for MNIST
Latency $1.47$ ms
Throughput $130.61$ imgs/ms
@@
6 Conclusion & Future Work
This work introduced a novel architecture combining MLPs and graph convolutional layers, benchmarked on popular grayscale image datasets. The model demonstrated strong performance and low complexity, highlighting the importance of lightweight, low-latency image classifiers for various applications. Its efficacy was shown across general image classification, SAR ATR, and medical image classification, with an FPGA implementation underscoring its hardware compatibility. Key innovations include using a single-layer GCN, which, along with batch-wise attention, enhances accuracy and reduces variance. Future work should explore extending this approach to color image datasets and other domains, optimizing the architecture for even greater efficiency, and further investigating the potential of graph neural networks in shallow models.

### v6

@@ Abstract
Image classifiers for domain-specific tasks like Synthetic Aperture Radar Automatic Target Recognition (SAR ATR) and chest X-ray classification often rely on convolutional neural networks (CNNs). These networks, while powerful, experience high latency due to the number of operations they perform, which can be problematic in real-time applications. Many image classification models are designed to work with both RGB and grayscale datasets, but classifiers that operate solely on grayscale images are less common. Grayscale image classification has critical applications in fields such as medical imaging and SAR ATR. In response, we present a novel grayscale image classification approach using a vectorized view of images. By leveraging the lightweight nature of Multi-Layer Perceptrons (MLPs), we treat images as vectors, simplifying the problem to grayscale image classification. Our approach incorporates a single graph convolutional layer in a batch-wise manner, enhancing accuracy and reducing performance variance. Additionally, we develop a customized accelerator on FPGA for our model, incorporating several optimizations to improve performance. Experimental results on benchmark grayscale image datasets demonstrate the effectiveness of our approach, achieving significantly lower latency (up to 16×16$\times$16 × less on MSTAR) and competitive or superior performance compared to state-of-the-art models for SAR ATR and medical image classification.
Index Terms—  GCN, grayscale, MLP, low-latency
@@ 1 Introduction
A particular instance where low-latency machine learning is needed is grayscale image classification for SAR ATR. For example, a targeting system on a satellite is costly, and decisions must be made using SAR efficiently and accurately. Examples like this are where low-latency grayscale image classification comes into play. It is often the case that image classifiers work on RGB datasets and grayscale image datasets, but seldom do modern image classifiers focus solely on the grayscale setting. RGB models are overkill for the grayscale setting, as the grayscale problem allows us to focus on a single channel. Models focusing on grayscale image classification are naturally more efficient, as they can concentrate on a single channel rather than three. Thus, many image classifiers that generalize to the grayscale image classification are not truly optimized for the grayscale case. For these reasons, we present a lightweight grayscale image classifier capable of achieving up to 16×16$\times$16 × lower latency than other state-of-the-art machine learning models on the MSTAR dataset.
From a trustworthy visual data processing perspective, the demand for grayscale image classification requires data to be collected from various domains with high resolution and correctness so that we can train a robust machine learning model. Additionally, recent advancements in machine learning rely on convolutional neural networks, which often suffer from high computation costs, large memory requirements, and many computations needed, resulting in poor inference latency, poor scalability, and weak trustworthiness.
The inherent novelties of our model are as follows: Our proposed method is the first to vectorize an image in a fully connected manner and input the resultant into a single-layer graph convolutional network (GCN). We also find that a single GCN layer is enough to stabilize the performance of our shallow model. Additionally, our proposed method benefits from a batch-wise attention term, allowing our shallow model to capture interdependencies between images and form connections for classification. Finally, by focusing on grayscale imagery, we can focus on a streamlined method for grayscale image classification rather than concentrating on the RGB setting. A result of these novelties is extremely low latency and high throughput for SAR ATR and medical image classification.
@@
  • Experiments show that our model achieves competitive or leading accuracy with respect to other popular state-of-the-art models while vastly reducing latency and model complexity for SAR ATR and medical image classification.
  • We implement our model on a state-of-the-art FPGA board, Xilinx Alveo U200. Compared with state-of-the-art GPU implementation, our FPGA implementation achieves comparable latency and throughput with state-of-the-art GPU, with only $1/41$ of the peak performance and $1/10$ of the memory bandwidth.
@@ 2 Problem Definition
Latency refers to the total time (from start to finish) it takes to gather predictions for a model in one batch. A lightweight machine learning model aims to maximize throughput and accuracy while minimizing latency.
@@ 4 Overview and Architecture
Our model vectorizes a batch of images, allowing us to use a fully connected layer (FC) pixel-wise for low computation time rather than relying on convolutional neural networks. We vectorize the input images and input them into a fully connected layer. Then, we use a graph convolutional layer to learn similarities between images batch-wise. We then apply a batch-wise attention term, which is inputted into an FC for classification. [^1] We make our code publicly available at https://github.com/GECCOProject/GECCO
@@
Fully Connected Layer. We input $\mathbf{X}_{1}$ into an FC layer with output dimensionality $D_{out}$. Formally, $\mathbf{X}_{2}=\sigma(\mathbf{X}_{1}\mathbf{W}_{1}+\mathbf{b}_{1})$, where $\sigma$ is the ReLU function, $\mathbf{W}_{1}$ is a learned weight matrix, $\mathbf{b}_{1}$ is a bias term, and $\mathbf{X}_{2}\in\mathbb{R}^{B\times D_{out}}$.
@@
Applying graph convolution to $\mathbf{X}_{3}$ yields $\mathbf{X}_{4}$. Formally, $\mathbf{X}_{4}=\sigma\left(\mathbf{A}\mathbf{X}_{3}\mathbf{W}_{2}\right)$, where $\mathbf{W}_{2}$ is a learned weight matrix and $\sigma$ represents the sigmoid function. After the graph convolution, we apply batch normalization and max-pooling operations to $\mathbf{X}_{4}$, resulting in a dimensionality of $\mathbb{R}^{B\times\lfloor D_{out}/2\rfloor}$.
@@
$\mathbf{X}_{5}$ $=\left(\frac{\sigma\left(\mathbf{X}_{4}\mathbf{X}^{\top}_{4}\right)}{\sum_{i=1}^{B}\sigma\left(\mathbf{X}_{4}\mathbf{X}_{4}^{\top}\right)_{i}}\right)\mathbf{X}_{4}$
@@
The residual connection is defined $\mathbf{X}_{6}=\mathbf{X}_{5}+\mathbf{X}_{4}$. The residual term makes the learning process easier and more stable. By multiplying a softmax-like term with the output of the previous graph convolution, we weigh the correspondence of each image compared to other similar images within similar images batch-wise. We then feed the residual term into an FC inputted into the softmax function for classification results.
@@ 5.1 Datasets
Datasets from several domains are examined to gauge the effectiveness of GECCO in diverse settings. We use the SAR ATR dataset, MSTAR, and a medical imaging dataset CXR [12].
@@
Our goal is to create a real-time system. That is, we wish to minimize the inference latency and maximize the throughput of our model while maintaining leading or competitive accuracy on its respective dataset. In the following sections, we measure the inference latency and throughput, as described in section 2.
@@ 5.2.1 Backbone
Table 1: Performance of GCN Layers on MSTAR
@@
GCN [13] $98.89\%$ $50.04\pm 6.85$ $5.86\pm 0.98$
TAGConv [14] $99.05\%$ $47.87\pm 7.11$ $6.24\pm 1.32$
SAGEConv [15] $99.08\%$ $51.77\pm 8.19$ $5.95\pm 0.87$
ChebConv [16] $98.56\%$ $45.37\pm 5.99$ $6.83\pm 1.27$
ResConv [17] $\bf{99.29\%}$ $\bf{52.98\pm 9.04}$ $\bf{5.22\pm 1.03}$
We use the following hyperparameters listed in Table 2 for our experiments.
@@ 5.2.2 Experimental Performance
We summarize our findings in Tables 3 and 4. We report the best-performing accuracy, average throughputs, and latencies with their standard deviations. Our model outperforms every other model in terms of throughput and latency across all datasets, leads accuracy on the MSTAR dataset, and performs competitively in terms of accuracy on all datasets.
@@
Swin-T $86.04\%$ $1.36\pm 0.10$ $46.98\pm 3.20$
SS-ViT $95.61\%$ $2.29\pm 0.43$ $27.97\pm 5.26$
VGG16 $93.13\%$ $1.69\pm 0.33$ $37.89\pm 7.52$
FastViT $91.78\%$ $1.04\pm 0.13$ $61.44\pm 7.69$
ResNet34 $98.64\%$ $3.13\pm 0.22$ $20.48\pm 1.39$
GECCO $\bf{99.29\%}$ $\bf{12.26\pm 2.42}$ $\bf{5.22\pm 1.03}$
@@
Swin-T $73.66\%$ $0.27\pm 0.05$ $236.71\pm 46.09$
SS-ViT $71.09\%$ $1.03\pm 0.21$ $62.35\pm 12.85$
VGG16 $\bf{82.01\%}$ $0.76\pm 0.25$ $84.10\pm 28.43$
FastViT $75.46\%$ $1.06\pm 0.14$ $60.30\pm 14.24$
ResNet34 $78.31\%$ $0.60\pm 0.11$ $105.84\pm 19.39$
GECCO $77.57\%$ $\bf{2.63\pm 0.55}$ $\bf{24.32\pm 5.08}$
@@ 5.2.3 Model Complexity Metrics
Model complexity metrics for this paper include the number of multiply-accumulate operations (MACs), the number of model parameters, the model size, and the number of layers. In other words, suppose accumulator $a$ counts an operation of arbitrary $b,c\in\mathbb{R}$. We count the number of multiply-accumulate operations as $a\leftarrow a+(b\times c)$. Additionally, the layer count metric is an essential factor of latency. Decreasing the number of layers will also improve the latency of a model’s inference time. The goal of an effective machine learning model is to maximize throughput while minimizing the number of MACs and the number of layers, in our case.
@@
GECCO $\bf{5.10\times 10^{4}}$ $\bf{5.08\times 10^{4}}$ $\bf{0.19}$ $\bf{16}$
@@ 5.2.4 Ablation Study
We perform an ablation study to verify that the components of our proposed model contribute positively to the overall accuracy of the MSTAR dataset.
@@
Mini-batch GNN Weighted Sum Residual Term Accuracy on MSTAR
$99.29\%$
$97.94\%$
$88.04\%$
$78.64\%$
@@
Fig. 2: Accuracy on MSTAR vs. Number of Graph Convolutional Layers
@@ 5.3 Discussion
It is difficult for our model to generalize to the RGB setting. We attribute this challenge to the vectorization process that our model uses. Learning on three channels poses a complexity challenge, as GECCO is very shallow and simple, thus making it challenging to learn on three separate channels. Additionally, our model is optimized for a low-complexity dataset regime, as datasets like CIFAR and ImageNet are much too complex for our shallow model, as they pose a complexity challenge.
Our proposed method does not make use of positional embeddings or class tokens. GECCO can learn essential features using the weighted residual term. Additionally, we tested the addition of positional embeddings and class tokens and found no improvement in accuracy across various datasets. We note that the $\mathbf{X}_{5}$ attention-like term adds positional awareness to the model.
@@ 5.4 FPGA Implementation
We develop an accelerator for the proposed model on state-of-the-art FPGA, Xilinx Alveo U200 [24], to further highlight the model’s efficiency and compatibility with hardware. It has 3 Super Logic Regions (SLRs), 4 DDR memory banks, 1182k Look-up tables, 6840 DSPs, 75.9 Mb of BRAM, and 270 Mb of URAM. The FPGA kernels are developed using the Xilinx High-level Synthesis (HLS) tool to expedite the design process.
@@
Our FPGA design incorporates several novel features: (1) Portability of the design: We design a parameterized hardware template using HLS. It is portable to different FPGA platforms, including embedded and data-center FPGAs. We present our hardware mapping algorithm in Algorithm 1. (2) Resource sharing: The model is executed layer-by-layer. Each layer in the model is decomposed into basic kernel functions. The basic kernel functions, including matrix multiplication, elementwise activation, column-wise and row-wise summations, max pooling, and various other elementwise operations, are implemented separately and subsequently invoked within their corresponding layers. Due to the reuse of these fundamental kernel functions across multiple layers, FPGA resources are shared among the different layers, maximizing runtime hardware resource utilization. (3) Single-load strategy: We employ a one-time data load strategy to load the required data from DDR only once. All other data required for the computations are stored in on-chip memory, reducing inference latency. Figure 3 illustrates the overall hardware architecture of our design.
@@
We utilize the Vitis tool [25] for hardware synthesis and place-and-route to determine the achieved frequency. The Vitis Analyzer tool is then used to report resource utilization and the number of clock cycles. The latency is calculated by multiplying the achieved clock period by the number of cycles. Table 8 reports the results obtained for the MSTAR dataset. Given the compact design and resource efficiency of the model, it can be accommodated within a single SLR. Hence, we deploy multiple accelerator instances across multiple SLRs, each with one instance. This increases the inference throughput. Table 8 shows the latency obtained for a single inference and the throughput achieved by running the design on 3 SLRs concurrently.
Table 7: Comparison with state-of-the-art GPU platform
GPU Our Design Platform NVIDIA A5000 Alveo U200 Technology Samsung 8 nm TSMC 16 nm Frequency 1.17 GHz 200 MHz Peak Performance (TFLOPS) 27.7 0.66 On-chip Memory 6 MB 35 MB Memory Bandwidth 768 GB/s 77 GB/s Latency on MSTAR (ms) 5.22 5.65 Throughput on MSTAR (imgs/ms) 12.26 33.98
Algorithm 1 Hardware Mapping Algorithm (See the definition of layer in Section 4)
1:Model $f()$ and the input images;
2:Execution result;
3:for each layer $i$ in $f()$ do
4:     if  layer $i$ is a fully connected layer  then
5:         Map to the matrix multiplication unit
6:     if layer $i$ is a graph convolution layer then
7:         Map to the matrix multiplication unit
8:         Map to the activation unit
9:         Map to the elementwise operation unit
10:         Map to the matrix addition unit
11:     if layer $i$ is a batch-wise attention then
12:         Map to the matrix multiplication unit
13:         Map to the activation unit
14:         Map to the elementwise operation unit
15:     if layer $i$ is a max pooling layer then
16:         Map to the max pooling unit
17:     if layer $i$ is a activation layer then
18:         Map to the activation unit
19:     if layer $i$ is a batch normalization layer  then
20:         Map to the batch normalization unit
@@
We compare our FPGA implementation with the baseline GPU implementation. The GPU baseline is executed on an NVIDIA RTX A5000 GPU, which operates at 1170 MHz and has a memory bandwidth of 768 GB/s. However, the FPGA operates at 200 MHz and has an external memory bandwidth of 77 GB/s. We compare the hardware features of the two platforms in Table 7. Although the GPU has higher peak performance (41×41$\times$41 ×) and memory bandwidth (10×10$\times$10 ×), our FPGA implementation achieves a comparable latency of $5.65~{}ms$ and an improved throughput of $33.98~{}imgs/ms$.
@@
Table 8: Resource Utilization (per SLR), Latency, and Throughput for MSTAR dataset
Latency $5.65$ ms
Throughput $33.98$ imgs/ms
@@
6 Conclusion and Future Work
This work introduced a novel architecture combining fully connected and graph convolutional layers, benchmarked on popular grayscale image datasets. The model demonstrated strong performance and low complexity, highlighting the importance of lightweight, low-latency image classifiers for various applications. Its efficacy was shown across SAR ATR and medical image classification, with an FPGA implementation underscoring its hardware friendliness. Key innovations include using a single-layer GCN, which, along with batch-wise attention, enhances accuracy and reduces variance. Future work should explore extending this approach to color image datasets and other domains, optimizing the architecture for even greater efficiency, and further investigating the potential of graph neural networks in shallow models.